OpenStudy (anonymous):

Brian, Chris, and Damien took a math test that had 20 questions. The number of questions Brian got right is 14 more than 1/4 the number of questions Chris got right. Damien correctly answered 2 less than 5/4 the number of questions Chris answered correctly. If Brian and Damien have the same score, which statement is true? Brian and Damien both answered 2 fewer questions correctly than Chris did Brian and Damien both answered 4 more questions correctly than Chris did Brian and Damien both answered 2 more questions correctly than Chris did Brian and Damien both answered 4 fewer questions

Directrix (directrix):

B represents the number of questions Brian got right C represents the number of questions Chris got right D represents the number of questions Damien got right

Directrix (directrix):

The number of questions Brian got right is 14 more than 1/4 the number of questions Chris got right. B = 14 + (1/4) * C

Directrix (directrix):

Damien correctly answered 2 less than 5/4 the number of questions Chris answered correctly. D = (5/4) C - 2

Directrix (directrix):

If Brian and Damien have the same score, 14 + (1/4) * C = (5/4) C - 2 Solve the equation for C.

Directrix (directrix):

@hhjhethhzd